OpenStudy (anonymous):

sqrt(1-x^2) + sqrt(1-y^2) = a(x-y) prove that dy/dx = sqrt[ (1-y^2)/(1-x^2) ]

OpenStudy (souvik):

put \[x=\sin \theta \] and \[y=\sin \phi \]....

OpenStudy (anonymous):

then we get \[\frac{\cos \theta + \cos \alpha }{ \sin \theta - \sin \alpha } = a\] then after that what to do next

OpenStudy (souvik):

well...use this \[\cos \theta + \cos \alpha =2 \cos(\theta+\alpha )/2*\cos(\theta -\alpha )/2\] and \[\sin \theta -\sin \alpha= 2\cos(\theta+\alpha )/2*\sin(\theta-\alpha )/2\]

OpenStudy (souvik):

do you know this formulas...?

OpenStudy (anonymous):

yes very much and thanks a lot

OpenStudy (souvik):

got your answers?